OpenStudy (thatonegirl_):

Find f'(x) for f(x)=4^x

OpenStudy (holsteremission):

Hint: \(4^x=e^{\ln 4^x}=e^{x\ln4}\).

OpenStudy (thatonegirl_):

ahh thank you